OpenStudy (aonz):

Please help some1 i need help Question below

OpenStudy (aonz):

If x + y = 1 and x^3 + x^3 = 19 find value of x^2 + y^2

OpenStudy (anonymous):

x^3 + x^3 = 19 ? 2x^3 = 19 x^3 = 19/2 x = cubed root of 19/2 u can substitute this into the first equation x + y = 1 to solve for y, then you can solve for x^2 + y^2

OpenStudy (aonz):

opps i meant x^3 + y^3 =19

OpenStudy (anonymous):

i thought thats what u meant thats why i put a ? after it :)

OpenStudy (aonz):

yes (a + b)(a^2 - ab + b^2)

OpenStudy (anonymous):

Great :) a³ + b³ = (a + b)(a² - ab + b²) 19 = 1 (a² - ab + b²)

OpenStudy (aonz):

umm ok... i still dont think i get the question

OpenStudy (anonymous):

i would just use substitution

OpenStudy (aonz):

how :P

OpenStudy (anonymous):

let y = 1 - x then y^3 = (1 - x)^3 expand it out then you can write x^3 + (1 - x)^3 = 19 should come out nicely

OpenStudy (anonymous):

a² + b² - ab = (a + b)² - 2ab -ab 19 = 1 - 3ab -> ab = -18/ 3 = -6